OpenStudy (anonymous):

Find the area of the region bounded by y=6-2x,y=x+6 and y=0. integrate with respect to y.

OpenStudy (ash2326):

@houseofaura can you graph the lines, y=6-2x, y=x+6 and find the intersection points.

OpenStudy (ash2326):

??

OpenStudy (anonymous):

(0,6)

OpenStudy (ash2326):

Good, I'll plot the lines

OpenStudy (ash2326):



OpenStudy (anonymous):

i need to find the area

OpenStudy (loser66):

divide the region into 2 parts: left and right from y -axis, take int respect to y

OpenStudy (anonymous):

i dont understand

OpenStudy (loser66):

since you ask to take int respect to y, if you don't do that, how?